# DESI Dark Energy Time Evolution is Recovered by Cosmologically Coupled Black Holes

### v1

@@ 2 Theory
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}+\frac{3}{a}\rho_{\Lambda}(1+w)=\frac{\Xi}{Ha^{4}}\psi\,.$ (2.3)
In the limit where $\Xi\to 0$, DE evolves solely through gravitation
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=-\frac{3}{a}\rho_{\Lambda}(1+w)\,.$ (2.4)
Given the $w_{0}w_{a}$ model in Eq. 1.1, this becomes
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=-\frac{3}{a}\rho_{\Lambda}\left[(1+w_{0}+w_{a})-w_{a}a\right]\,.$ (2.5)
This equation is separable, and the solution is
$\rho_{\Lambda}=N\frac{e^{3aw_{a}}}{a^{3(1+w_{0}+w_{a})}}\,,$ (2.6)
where $N$ is a normalization constant. Note that, for $w_{0}+w_{a}<-1$, the solution can become “phantom,” growing from zero without any source. To reproduce the present-day DE density,
$N=\rho_{\mathrm{cr}}\Omega_{\Lambda}e^{-3w_{a}}\,,$ (2.7)
assuming Robertson-Walker (RW) flatness $\Omega_{\mathrm{K}}:=0$ to write
$\Omega_{\Lambda}=1-\Omega_{m}$ (2.8)
to a good approximation. The evolution of DE density under the $w_{0}w_{a}$ model, described by $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ in Eq. 2.5 can be expressed entirely by $a$ with Eq. 2.6. This derivative, for all DESI+CMB+SNIa most-likely $w_{0}$ and $w_{a}$ is shown in solid lines of Figure 1. The BH DE model is characterized by $w:=-1$ and $\Xi\neq 0$. In this setting, Eq. 2.3 becomes
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=\frac{\Xi}{Ha^{4}}\psi\,.$ (2.9)
For simplicity, we approximate the Hubble expansion rate $H$ as follows: since $\psi\to 0$ as $z\to 0$, then $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a\sim 0$ for $z\lesssim 1$. For $z\gtrsim 1$, the universe is matter-dominated and the impact of $\rho_{\Lambda}$ is small. Provided that baryon consumption is not too large, the typical expression for the Hubble rate
$H=H_{0}\sqrt{\Omega_{m}/a^{3}+(1-\Omega_{m})}\,,$ (2.10)
remains approximately valid. Substitution into Eq. 2.9 gives
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=\frac{\Xi\psi(a)a^{-5/2}}{H_{0}\sqrt{\Omega_{m}+(1-\Omega_{m})a^{3}}}\,.$ (2.11)
This derivative is shown for all DESI+CMB+SNIa best-fit $\Omega_{m}$ and $H_{0}$ in the dotted lines of Figure 1. To ensure the correct DE density today is recovered
$\Xi=\rho_{\mathrm{cr}}\Omega_{\Lambda}\left[\int_{a_{i}}^{1}\frac{\psi(a)a^{-5/2}}{H_{0}\sqrt{\Omega_{m}+(1-\Omega_{m})a^{3}}}~{}\mathrm{d}a\right]^{-1}\,.$ (2.12)
The value of $\Xi$ for each DESI+CMB+SNIa measurement of $\Omega_{m}$ and $H_{0}$ is provided in the Figure 1 legends.
Figure 2: Impact of different SFRDs on the $w_{0}w_{a}$ model fit to the rate of production of dark energy. Shaded regions indicates the DESI+CMB+DESY5 $w_{0}w_{a}$ measurement with central value $w_{0}$. The $w_{a}$ values are taken uniformly within twice the $68\%$ confidence regions. The displayed $w_{0}w_{a}$ fits to the BH DE model with different SFRDs span a reasonable range in uncertainty. The Harikane SFRD [84] is a direct measurement, from rest-frame UV bright sources found by JWST, but may underestimate the total SFRD. The Trinca SFRD [85] is an estimate of the total SFRD from a semi-analytic model.
@@ 3 Results
We can now answer the question: If the production of DE is described by cosmologically coupled BH production, what conclusions can be drawn about DE evolution when the data are interpreted using a $w_{0}w_{a}$ model? To investigate this, we perform a least-squares fit of the $w_{0}w_{a}$ model to the BH DE model defined by Eqs. 2.11 and 2.12. We consider each DESI+CMB+SNIa measurement for $\Omega_{m}$ and $H_{0}$ and adopt the standard Madau and Fragos parametrization for $\psi$ [83]. The results, shown in Figure 1 with dashed lines, indicate a good match to the DESI measured solid lines across all three cases. The least-squares $w_{0}w_{a}$ fit to BH DE falls entirely within $1\sigma$ of the DESI most-likely $w_{0}$ and $w_{a}$ for all data sets. The normalization of the BH DE model is determined by $\Omega_{m}$ and $H_{0}$, yielding $\Xi\simeq 5$ in all three scenarios.
Next, we examine the sensitivity of our results to the SFRD. While the SFRD at $z\lesssim 4$ is relatively well-constrained, the SFRD at higher redshifts remains uncertain. In Figure 2, we compare the results using the Madau and Fragos SFRD model across all redshifts to other SFRD models informed from James Webb Space Telescope (JWST) results at high redshifts. The Harikane SFRD [84], based on UV-bright measurements, is expected to provide a lower-bound on SFRD at $z>5$ (see also [86]). The Trinca SFRD, a $z>4$ prediction from the semi-analytic Cosmic Archaeology Tool, is informed by recent JWST data and suggests a slower initial decline in SFRD as redshift increases[85, 87]. For this SFRD at $z\leqslant 4$, we use the Madau and Fragos SFRD scaled by 2$\sim$ 2$\times$ 2 × for continuity at the junction. For simplicity, we present only the DESI+CMB+DESY5 data set for comparison, adopting the central $w_{0}$ and twice the $68\%$ confidence region for $w_{a}$. The other SFRDs show consistency at $\sim 1\sigma$ and $\sim 2\sigma$ levels respectively, leading us to conclude that our results align with contemporary estimates of the high redshift SFRD. We note that DE production at $z>10$ is not well-parameterized by the $w_{0}w_{a}$ model, which must adopt a prior so that $\rho_{\Lambda}\to 0$ as $a\to 0$.
@@ 4 Discussion
Physically, $\Xi\sim 5$ implies that, on average, $5M_{\odot}$ of baryons convert into BH DE for every $1M_{\odot}$ of stellar material produced. From Eq. 2.1, the fraction of baryons remaining today, relative to stellar first light, is approximately:
$\rho_{b}(a=1)/\Omega_{b}\rho_{\mathrm{cr}}\sim 0.54\,.$ (4.1)
We have used $\Omega_{b}h^{2}=0.02205$ [88, Eqn. (24.6)], with $h^{2}=(H_{0}/100)^{2}$, where $H_{0}$ is measured in units of $\mathrm{km}~{}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ depends on the DESI dataset adopted. This result implies that approximately $\sim 45\%$ of baryons are missing at late times.
In fact, there has been a “missing baryons problem” for over a decade, with measurements indicating a shortfall of about $30\%$ [89]. Recent estimates pioneering fast-radio bursts (FRBs) report a value $\Omega_{b}=0.05^{+0.021}_{-0.025}$ at $95\%$ confidence level [90],555c.f. text below their Eqn. 8 in Methods, where their central value drops by $20\%$ at comparable uncertainties although direct constraints still indicate a shortfall of around $\sim 75\%$ [91, Fig. 1]. Our estimated survival fraction is marginally consistent with the FRB measurement, but we believe the actual consumption is likely lower. At face value, $\sim 45\%$ consumption would result in approximately a $\sim 10\%$ drop in $\Omega_{m}$ across $41<z<$ 4, enough to shift the onset of accelerated expansion earlier within DESI’s sensitivity. For consistency with DESI-1YR’s reported results, we do not incorporate this effect into Eq. 2.11, leaving it for future study.
@@
This consumption of baryons may help explain other unexpected aspects of the DESI-1YR analysis. DESI-1YR results find a most probable summed neutrino mass $\sum m_{\nu}$ equal to zero [Fig. 11, Left Panel 8](see also [92]), contradicting the computed $\sum m>0.059~{}\mathrm{eV}$ based on $\Delta m_{ij}^{2}$ constraints from neutrino oscillation measurements [e.g. 93, 94]. [^6] This standard constraint follows from algebra $\sqrt{\Delta m_{21}^{2}}+\sqrt{\Delta m_{31}^{2}}=(m_{2}-m_{1})+(m_{3}-m_{1})<m_{2}+m_{3}<\sum m_{\nu},$ (4.2) under the assumption that $m_{\nu}\geqslant 0~{}\forall\nu$ and $m_{i}$ are monotonically increasing in $i$ (normal hierarchy). The summed neutrino mass approaching zero in DESI-1YR can be interpreted as a preference for lower matter density at late times, relative to that inferred from the CMB. [^7] It has also been proposed that a preferred zero summed neutrino mass could arise from new physics in the neutrino sector [95] We suspect that the baryon consumption required in the BH DE scenario will allow $\sum m_{\nu}$ to increase towards physically reasonable values. It is important to note that $\Omega_{b}h^{2}$ and $\sum m_{\nu}$ are not entirely degenerate, as the required DE density constrains the amount of baryon consumption.

### v2

@@ 2 Theory
$\frac{\mathrm{d}\rho_{\mathrm{DE}}}{\mathrm{d}a}+\frac{3}{a}\rho_{\mathrm{DE}}(1+w)=\frac{\Xi}{Ha^{4}}\psi\,.$ (2.4)
The BH DE model is characterized by $w:=-1$ and $\Xi\neq 0$. In this setting, Eq. 2.4 becomes
$\frac{\mathrm{d}\rho_{\mathrm{DE}}}{\mathrm{d}a}=\frac{\Xi}{Ha^{4}}\psi\,.$ (2.5)
The Hubble expansion rate $H$ is determined algebraically in terms of the physical densities $\rho_{i}$ via the Friedmann energy equation,
$H^{2}=\left(\frac{8\pi G}{3}\right)\left[\rho_{b}(a)+\rho_{\gamma}(a)+\rho_{\nu}(a)+\rho_{\mathrm{DE}}(a)+\rho_{c}(a)\right],$ (2.6)
where $\gamma$, $\nu$ and $c$ denote typical photon, neutrino, and cold dark matter contributions respectively [e.g. 90]. We omit a curvature contribution $\propto 1/a^{2}$, so the universe is always spatially flat and critical. Combined with the derivative of Eq. 2.2 with respect to $a$, Eqs. 2.5 and 2.6 give a closed system of differential equations for $a\geqslant a_{i}$
$\begin{split}\frac{\mathrm{d}\rho_{\mathrm{DE}}}{\mathrm{d}a}&=\frac{\Xi}{Ha^{4}}\psi\,\\ \frac{\mathrm{d}\rho_{b}}{\mathrm{d}a}&=-\left(\frac{3\rho_{b}}{a}+\frac{\mathrm{d}\rho_{\mathrm{DE}}}{\mathrm{d}a}\right)\\ \frac{\mathrm{d}D_{\mathrm{M}}}{\mathrm{d}a}&=\frac{1}{Ha^{2}}.\end{split}.$ (2.7)
We include the comoving distance $D_{\mathrm{M}}$ [e.g. 91], as it is required to interpret DESI BAO. It can be determined by adjusting the integration limits (and sign) using the Fundamental Theorem of Calculus afterward so that $D_{\mathrm{M}}(1):=0$. Note that for flat Robertson-Walker cosmologies, the transverse comoving distance is equal to the line-of-sight comoving distance. Initial conditions for all species are determined by projecting species’ $\omega_{j}^{\mathrm{proj}}$ values to $a_{i}$. We define the DE density $\rho_{\mathrm{DE}}(a\leqslant a_{i}):=0$. For $\rho_{\nu}$, we follow DESI [12] and adopt $N_{\mathrm{eff}}:=3.044$, a single massive neutrino species with $m_{\nu}:=0.06~{}\mathrm{eV}$, and track their energy density exactly [e.g. 92, §3.3].
3 Methods
Model / Parameter dynesty Prior dynesty Best-Fit dynesty Posterior Deviation
CCBH Trinca $\bm{\psi}$ $\left(\chi^{2}=12.66\right)$
$\Xi$ $\mathcal{U}[0,10]$ $1.403$ $1.396^{+0.050}_{-0.048}$ $0.15\sigma$
$\omega_{c}$ $\mathcal{U}[0.01,0.4]$ $0.1237$ $0.1240^{+0.0083}_{-0.0078}$ $-0.04\sigma$
$100\omega_{b}^{\mathrm{proj}}$ $\mathcal{N}(2.218,0.055)$ $2.238$ $2.219\pm 0.054$ $0.35\sigma$
Flat $\bm{\Lambda}$CDM $\left(\chi^{2}=12.74\right)$
$H_{0}~{}[\,{\rm km\,s^{-1}\,Mpc^{-1}}]$ $\mathcal{U}[20,100]$ $67.72$ $68.71^{+0.81}_{-0.80}$ $-1.23\sigma$
$\omega_{c}$ $\mathcal{U}[0.01,0.4]$ $0.1129$ $0.1163^{+0.0082}_{-0.0081}$ $-0.42\sigma$
$100\omega_{b}^{\mathrm{proj}}$ $\mathcal{N}(2.218,0.055)$ $2.101$ $2.219\pm 0.055$ $-2.16\sigma$
Table 1: Fit configuration and results for primary dynesty analysis of DESI Year 1 BAO data. Best-fit parameter values, with $\chi^{2}$, and posteriors distributions at $68\%$ confidence are given. Priors are either uniform $\mathcal{U}$ on the indicated range or normally distributed $\mathcal{N}(\mu,\sigma)$. Projected comoving baryon density $\omega_{b}^{\mathrm{proj}}$ has been multiplied by $100$ for clarity, and is taken from Big Bang Nucleosynthesis (BBN) constraints. For the CCBH model, the present-day comoving baryon density is less than $\omega_{b}^{\mathrm{proj}}$ due to baryon conversion into DE inside BHs. Deviation gives the discrepancy in best-fit values relative to posterior maxima, in units of $\sigma$. Note that $\Lambda$CDM best-fit $\omega_{b}^{\mathrm{proj}}$ pulls away from the posterior mean, which is dominated by the BBN prior. This indicates a preference in the BAO data for lower $\omega_{b}^{\mathrm{proj}}$.
Figure 2: Deviations of best-fit models, relative to Planck $\Lambda$CDM, in comoving distance $D_{\mathrm{M}}$ (top), Hubble distance $D_{H}$ (middle), and a volume-averaged combination $D_{V}$ (bottom). DESI BAO datapoints, relative to Planck $\Lambda$CDM, are displayed with $68\%$ confidence (black dots). When CCBHs provide the physical source of DE (solid, purple), the resulting expansion history closely tracks a $\Lambda$CDM model (dashed, orange) for $z<2.5$, but this model is not equal to the $\Lambda$CDM model inferred from early-universe experiments like Planck (dotted, green). DESI reports that these two $\Lambda$CDM models are discrepant at $1.9\sigma$, qualitatively consistent with a naive $\chi^{2}\sim 21$ between the Planck best-fit model curve and the 12 DESI BAO degrees of freedom. Increased statistics from DESI Y3, Y5, and Y7 datasets, combined with additional constraints from Redshift Space Distortions (not present in the DESI Y1 initial release), should significantly tighten uncertainties and may adjust central values.
Figure 3: Posterior distributions for DESI BAO+BBN fits to DE sourced by cosmologically coupled BHs (purple), and $\Lambda$CDM (orange). Planck posteriors are shown (green) for comparison. Both fits draw the comoving baryon density inferred from early-universe measurements $\omega_{b}^{\mathrm{proj}}$ from a Gaussian BBN prior, which dominates the posterior, and has been omitted for clarity. Both fits draw the comoving cold dark matter density $\omega_{c}$ from the same uninformative prior. For CCBH, we draw $\Xi$: the amount of baryonic matter becoming BHs, per unit stellar material; while for $\Lambda$CDM, we draw $H_{0}$. For CCBH, the baryon survival fraction $s$ is also displayed. The posterior for $s$ implies that $\sim 30\%$ of baryons are genuinely lost, having been converted into DE inside BHs. This present-day baryon abundance is consistent with all contemporary baryon censuses.
Our goal is to determine the DESI BAO best-fit background-only CCBH and $\Lambda$CDM cosmologies. By fitting $\Lambda$CDM, we validate our independent pipeline against that of DESI. For CCBH, we fit for the collapse fraction $\Xi$, the free parameter that replaces $H_{0}$ used by DESI when fitting background-only $\Lambda$CDM with a BBN prior. For our primary analysis, we adopt a redshift $z>4$ SFRD that includes contributions from intrinsically faint objects [5, Fig. 3], while accurately reproducing $M_{UV}<-17$ observations from JWST [e.g. 93, Fig. 8] (although [c.f. 94, 95]). For $z\leqslant 4$, we adopt the standard Madau & Dickinson [6, 7] SFRD, enforcing continuity by scaling the normalization to match at $z=4$. The required factor $\sim 2$ is consistent with radio astrometric measures of the SFRD from synchrotron emission in protostellar disks [96]. These measurements are free from dust-obscuration systematics that impact all IR, UV, and optical inference of the SFRD and likely better determine the normalization. This combination of $z>4$ and $z<4$ behavior we will designate as “Trinca $\psi$.” We additionally compute with the standard Madau & Dickinson SFRD, designated as “Madau $\psi$,” to provide a fiducial visual reference in Figure 1.555Empirical determinations of the SFRD necessarily assume a particular cosmology to convert observed luminosities into stellar production rates, and to convert object counts into comoving densities. Because the expansion history does not begin to deviate from $\Lambda$CDM until $z\sim 1$ when star-formation is winding down, and the corrections are small relative to the original data uncertainties, we have neglected this correction.
Our code does not track recombination thermodynamics, so we adopt [97, Eqn. (16)]
$r_{\mathrm{d}}=55.154\frac{\exp\left[-72.3\left(\omega_{\nu}^{\mathrm{proj}}+0.0006\right)^{2}\right]}{\left(\omega_{c}^{\mathrm{proj}}+\omega_{b}^{\mathrm{proj}}\right)^{0.25351}\left(\omega_{b}^{\mathrm{proj}}\right)^{0.12807}}~{}\mathrm{Mpc}\,,$ (3.1)
to compute the baryon drag scale $r_{\mathrm{d}}$ to better than $0.06\%$ [^6] This model for $r_{\mathrm{d}}$ assumes $N_{\mathrm{eff}}=3.046$ and delivers accuracy to $0.021\%$. This $N_{\mathrm{eff}}$ is larger than our adopted $N_{\mathrm{eff}}=3.044$ by $0.06\%$, so we report the larger uncertainty. Next, define:
$D_{H}(z):=\frac{c}{H(z)}\qquad\qquad D_{V}(z):=\left[zD_{\mathrm{M}}(z)^{2}D_{H}(z)\right]^{1/3}.$ (3.2)
DESI can extract a BAO scale, relative to $r_{\mathrm{d}}$, in three ways: 1) fixed redshift and variable sky position via the transverse comoving distance $D_{\mathrm{M}}/r_{\mathrm{d}}$; 2) fixed sky position and variable redshift via the Hubble distance $D_{H}/r_{\mathrm{d}}$; and 3) the “angle averaged” mixture of these two signals $D_{V}/r_{\mathrm{d}}$ when statistics are not sufficient to report $D_{\mathrm{M}}/r_{\mathrm{d}}$ and $D_{H}/r_{\mathrm{d}}$ separately. The DESI-1YR BAO data include five distinct measurements each of $D_{\mathrm{M}}/r_{\mathrm{d}}$, $D_{H}/r_{\mathrm{d}}$, and two of $D_{V}/r_{\mathrm{d}}$ (see Figure 2). The best-fit central values $\mathbf{\mu}$ and covariance matrix for these 12 data points $\Sigma$ is provided by the DESI collaboration. [^7] This data can be found at https://github.com/CobayaSampler/bao_data/tree/master, and is significantly more precise than the data as listed in DESI Table 1. We solve Eq. 2.7 with scipy.odeint to determine model-predicted values $\mathbf{X}$ for the these 12 data points at their respective redshifts of measurement $z_{\mathrm{eff}}$. Because the variables are Gaussian distributed,
$\log\mathcal{L}_{\mathrm{DESI}}(\mathbf{X}):=-\frac{1}{2}(\mathbf{X}-\mathbf{\mu})^{\top}\Sigma^{-1}(\mathbf{X}-\mathbf{\mu}),$ (3.3)
where $\top$ denotes transpose and $\Sigma^{-1}$ denotes the inverse covariance matrix. We additionally enforce a hard cutoff in late-time baryon density based on confirmed baryon census in the local universe [98],
$\log\mathcal{L}_{b}\left(s\right):=\begin{cases}0&s>0.25\\ -\infty&s\leqslant 0.25\end{cases},$ (3.4)
where $s$ is the baryon survival fraction,
$s:=\frac{\rho_{b}(1)}{\rho_{b}(a_{i})a_{i}^{3}}.$ (3.5)
The combined $\log$ likelihood becomes
$\log\mathcal{L}:=\log\mathcal{L}_{\mathrm{DESI}}+\log\mathcal{L}_{b}.$ (3.6)
For consistency, we assert the same priors for CCBHs and $\Lambda$CDM: the DESI-adopted BBN prior on $\omega_{b}^{\mathrm{proj}}$ and a uniform prior in $\omega_{c}$. These are described in Table 1. Although DESI fits using $\Omega_{m}$, our $\Lambda$CDM posterior distributions agree with DESI to better than $0.5\%$ precision. [^8] We verified our pipeline to $<0.02\%$ precision against astropy, which is 25×25$\times$25 × smaller than $0.5\%$. To investigate this discrepancy, we constructed a secondary pipeline by modifying a version of CLASS [99]. A MontePython [100, 101] fit using an independently developed DESI Y1 BAO likelihood recovered the same values as our primary pipeline to $<0.02\%$ precision. We conclude our inference to be robust. In $\Lambda$CDM, the (fixed) DE density is determined by flatness and choice of $H_{0}$. For CCBHs, the DE density evolves away from zero according to Eq. 2.7. Thus, at our background-order study, CCBHs (given some adopted $\psi$) and $\Lambda$CDM have the same number of free parameters, but CCBHs give time-evolution with two fewer parameters than $w_{0}w_{a}$ models.
We use the dynamic nested sampling framework dynesty [102, 103, 104, 105, 106] [^9] The version we used can be found at doi.org/10.5281/zenodo.3348367, to measure the Bayesian evidence for Eq. 3.6 and estimate posterior distributions for $\omega_{b}^{\mathrm{proj}}$, $\omega_{c}$, $H_{0}$ and $\Xi$ for each cosmology. With these distributions, we compute derived distributions for $H_{0}$, $\Omega_{m}$, and the survival fraction $s$. The survival fraction for $\Lambda$CDM is expected to be unity to integrator precision, and serves as a further internal consistency check on our primary pipeline. We find $s_{\Lambda\mathrm{CDM}}=1.00015$, in agreement with expectations.
@@ 4 Results
Figure 1 shows the evolution of DE density $\rho_{\mathrm{DE}}$ for CCBH cosmologies, adopting the two SFRD models described in Section 3. Overlaid in color are the ±$\pm 1\sigma$ regions implied by DESI $w_{0}w_{a}$ parameters for each of the three Planck+DESI+SN data sets. Note that the SNe regions have the “hump” shape and the CCBH curves have the asymptoting “plateau” shape, as expected from monotonically decreasing star-formation for $z<2$. Table 1 summarizes all best-fit parameter values, $\chi^{2}:=-2\log\mathcal{L}$ for the best-fits, and posterior means with $68\%$ confidence regions. For CCBH Trinca $\psi$, the best-fit parameters all lie within ±$\pm 1\sigma$ of their posterior means. For flat $\Lambda$CDM, the best-fit $H_{0}$ and $\omega_{b}$ pull away from the posterior mean at $<-1\sigma$ and $<-2\sigma$, respectively. Because the posterior is the prior-weighted likelihood by Bayes’ Theorem, the discrepancy in $\omega_{b}$ indicates that the data prefer a lower comoving baryon density than Flat $\Lambda$CDM can accommodate.
Figure 2 shows our determined distance measures $D_{\mathrm{M}}/r_{d}$, $D_{H}/r_{d}$, and $D_{V}/r_{d}$, with Flat $\Lambda$CDM (Planck) distance measures subtracted off. It is not yet possible to distinguish late-time Flat $\Lambda$CDM (DESI 1YR+BBN) from CCBH Trinca $\psi$ (DESI 1YR+BBN) at background-order. DESI reports that their (late-time) $\Lambda$CDM is, however, distinct from Planck (early-time) $\Lambda$CDM at $1.9\sigma$. If this discrepancy should increase, then the $\Lambda$CDM model will be unable to reconcile early and late universe cosmological probes. However, CCBH can provide the missing late-time physics to reconcile these two datasets. Improved statistics from DESI data, and a comprehensive first-order study of CCBH with the Planck likelihoods, will help clarify whether or not this is the case.
Figure 3 summarizes the posterior distributions from our primary CCBH Trinca $\psi$ and Flat $\Lambda$CDM analyses. For comparison, we also include Planck TT,TE,EE+lowE+lensing posteriors. We have omitted $\omega_{b}^{\mathrm{proj}}$ because the posterior distribution for $\omega_{b}^{\mathrm{proj}}$ tracks the prior (see Table 1). The present-day expansion rate $H_{0}=(69.94\pm 0.81)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ is notably higher than both Flat $\Lambda$CDM (DESI 1YR+BBN) and the Planck value, because the consumption of baryons allows earlier transition to DE domination. We highlight that this value agrees to $0.5\%$ with $H_{0}=(69.59\pm 1.58)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ reported by the Chicago-Carnegie Hubble Program using Cepheids, the Tip of the Red Giant Branch (TRGB), and the J-Region Asymptotic Giant Branch stellar distance-ladder calibrations [107]. In the CCBH Trinca $\psi$ cosmology, the recovered value for $\omega_{c}$ is higher than Planck, allowing this CCBH cosmology to accommodate more background matter density contribution from massive neutrinos. In contrast, the Flat $\Lambda$CDM cosmology requires a lower $\omega_{c}$ than Planck in order to achieve the smaller $\Omega_{m}$ preferred by DESI BAO.
@@ 5 Discussion
We have reproduced the BAO data measured by DESI with a model that does not require adjustments to early-universe physics, and that leverages a well-motivated astrophysical source for DE: stellar-collapse CCBHs. This is the only known model that explains the coincidence problem of why DE has become relevant now: because it didn’t exist until stars formed. This also explains why the DE density is of the same order of magnitude as the matter density: DE is sourced by matter. Physically, $\Xi=1.39$ implies that, on average, $1.39M_{\odot}$ of baryons convert into BH DE for every $1M_{\odot}$ of stellar material produced. This value of $\Xi$ is recovered when $s=70\%$, implying that approximately $30\%$ of baryons are missing at late times. In the CCBH model, this means that those baryons are truly lost, being converted into DE.
@@
In fact, there has been a well-known “missing baryons problem” for over a decade, with measurements indicating a shortfall of about $30\%$ [108]. Our estimated survival fraction is consistent with all contemporary baryon censuses [98, Fig. 1]. This consumption of baryons may help explain other unexpected aspects of the DESI-1YR analysis. DESI-1YR results find a most probable summed neutrino mass $\sum m_{\nu}$ equal to zero [Fig. 11, Left Panel 12](see also [109]), contradicting the computed $\sum m_{\nu}>0.059~{}\mathrm{eV}$ based on $\Delta m_{ij}^{2}$ constraints from neutrino oscillation measurements [e.g. 110, 111]. [^10] This standard constraint follows from algebra $\sqrt{\Delta m_{21}^{2}}+\sqrt{\Delta m_{31}^{2}}=(m_{2}-m_{1})+(m_{3}-m_{1})<m_{2}+m_{3}<\sum m_{\nu},$ (5.1) under the assumption that $m_{\nu}\geqslant 0~{}\forall\nu$ and $m_{i}$ are monotonically increasing in $i$ (normal hierarchy). The summed neutrino mass approaching zero in DESI-1YR can be interpreted as a preference for lower matter density at late times, relative to that inferred from the CMB. [^11] It has also been proposed that a preferred zero summed neutrino mass could arise from new physics in the neutrino sector [112] We suspect that the baryon consumption required in the BH DE scenario will allow $\sum m_{\nu}$ to increase towards physically reasonable values. At background order, $\Omega_{b}h^{2}$ and $\sum m_{\nu}$ are not entirely degenerate in a CCBH cosmology, as the required DE density constrains the amount of baryon consumption. We note that $\sum m_{\nu}$ is also sensitive to the growth function at first-order in cosmological perturbations, which can be altered by novel CCBH dynamics [e.g. 64, §4]. We further expect that the decreased baryon density along the line of sight will slightly increase $z_{\mathrm{reio}}$, because $\tau_{\mathrm{reio}}$ is predominantly determined by the low $\ell$ multipoles of the CMB EE polarization power-spectrum. A comprehensive investigation of these, and other first-order observables, is the subject of future work.